# HoTPP Benchmark: Are We Good at the Long Horizon Events Forecasting?

### v1

@@ Abstract
In sequential event prediction, which finds applications in finance, retail, social networks, and healthcare, a crucial task is forecasting multiple future events within a specified time horizon. Traditionally, this has been addressed through autoregressive generation using next-event prediction models, such as Marked Temporal Point Processes. However, autoregressive methods use their own output for future predictions, potentially reducing quality as the prediction horizon extends. In this paper, we challenge traditional approaches by introducing a novel benchmark, HoTPP, specifically designed to evaluate a model’s ability to predict event sequences over a horizon. This benchmark features a new metric inspired by object detection in computer vision, addressing the limitations of existing metrics in assessing models with imprecise time-step predictions. Our evaluations on established datasets employing various models demonstrate that high accuracy in next-event prediction does not necessarily translate to superior horizon prediction, and vice versa. HoTPP aims to serve as a valuable tool for developing more robust event sequence prediction methods, ultimately paving the way for further advancements in the field. The code is available at https://github.com/ivan-chai/hotpp-benchmark.
@@ 1 Introduction
The world is full of events. Internet activity, e-commerce transactions, retail operations, clinical visits, and numerous other aspects of our lives generate vast amounts of data in the form of timestamps and related information. In the era of AI, it is crucial to develop methods capable of handling these complex data streams. We refer to this type of data as Event Sequences (ESs). Event sequences differ fundamentally from other data types. Unlike tabular data [wang2022transtab], ESs include timestamps and possess an inherent order. In contrast to time series data [lim2021timeseriessurvey], ESs are characterized by irregular time intervals and structured descriptions of each event. These differences necessitate the development of specialized models and evaluation practices.
The primary task in the domain of ESs is sequence modeling, specifically predicting the next event type and its occurrence time. Indeed, the ability to accurately forecast sequential events is vital for applications such as stock price prediction, personalized recommendation systems, and early disease detection. The simplified domain, represented as pairs of event types and times, is typically called Marked Temporal Point Processes (MTPP) [rizoiu2017hawkes-tutorial]. Additionally, structured modeling of dependencies between different data fields [mcdermott2024eventstreamgpt] can be considered an extension of MTPP.
In practice, a common question arises: what events will occur, and when, within a specific time horizon? Forecasting multiple future events poses unique challenges distinct from traditional next-event prediction tasks. The conventional approach to this problem typically involves the autoregressive application of next-event prediction models. While these models have demonstrated effectiveness in predicting the immediate next event, their performance tends to decline as the prediction horizon extends. This results in increased computational costs and reduced predictive accuracy over longer sequences. Despite this, it is commonly assumed that the quality of next-event prediction directly influences the quality of the corresponding autoregressive model.
In this paper, we challenge this belief and demonstrate that autoregressive models can be suboptimal for long-horizon predictions in some cases. To address this, we introduce HoTPP (Horizon Temporal Point Process), a novel benchmark designed to evaluate the effectiveness of event sequence prediction models over horizons. Our benchmark leverages a new metric, T-mAP (Temporal mean Average Precision), inspired by object detection methodologies in computer vision. We demonstrate that T-mAP offers a more nuanced assessment of a model’s performance in scenarios where precise time step predictions are challenging. This innovation addresses the inherent limitations of traditional metrics, which often fail to capture the complexities involved in long-term event forecasting.
Thus, our main contributions can be summarized as follows:
@@
    We introduce HoTPP, a new open-source benchmark designed to facilitate research on long horizon Event Sequences (ESs) prediction. This benchmark integrates datasets and methods from various domains, including financial transactions, social networks, medicine, and recommender systems.
@@
    We demonstrate that widely used evaluation methods for ESs and MTPPs often overlook critical aspects of model performance. Notably, we show that simple statistical baselines can sometimes outperform popular methods in terms of Optimal Transport Distance (OTD). To address the shortcomings of OTD, we introduce a new evaluation metric for MTPP called Temporal mAP (T-mAP) (Fig. 1).
@@
    We show that the length of the predicted sequence significantly impacts long-horizon prediction performance and must be carefully tuned. We also provide a method for selecting this hyperparameter.
@@
    We compare popular approaches from different domains and show that intensity-based MTPP modeling methods can be suboptimal compared to simple approaches. We also show that autoregressive approaches can be outperformed by direct prediction of multiple future events in some cases.
@@
Figure 1: Temporal mAP (T-mAP) evaluation pipeline. Unlike previous methods, T-mAP evaluates sequences of variable lengths over the prediction horizon. It also enhances performance assessment by analyzing label distributions rather than fixed predictions.
2 Background
2.1 Marked Temporal Point Processes
@@
Marked Temporal Point Process (MTPP) is a stochastic process that consists of a sequence of pairs $(t_{1},l_{1}),(t_{2},l_{2}),\dots$, where $t_{1}<t_{2}<\dots$ represent event times and $l_{i}\in\{1,\dots,L\}$ are event type labels. Common MTPP modeling approaches primarily focus on predicting the next event. A straightforward solution is to predict time and type independently. A more advanced approach involves splitting the original sequence into subsequences, one for each event type, and then independently modeling each subsequence’s time. In both approaches, the fundamental problem is modeling the arrival time of the next event. The stochastic process that generates these times is typically called a Temporal Point Process [rizoiu2017hawkes-tutorial]. For the details on MTPP modeling, refer to Appendix A.
2.2 Evaluation
MTPPs are typically evaluated based on the quality of next-event predictions. Time and type predictions are usually assessed independently. The quality of type prediction is characterized by the error rate, while time prediction error is measured using either the Mean Absolute Error (MAE) or Root Mean Squared Error (RMSE). Recent studies have advanced the measurement of horizon prediction quality by employing the Optimal Transport Distance (OTD). The OTD metric is analogous to the edit distance between the predicted event sequence and the ground truth. Below, we provide the formal definition of OTD.
@@ 2.2 Evaluation
Suppose there is a predicted sequence $S_{p}=\{(t^{p}_{i},l^{p}_{i})\}_{i=1}^{n_{p}},0<t_{1}^{p}<t_{2}^{p}<\dots$, and a ground truth sequence $S_{gt}=\{(t^{gt}_{j},l^{gt}_{j})\}_{j=1}^{n_{gt}},0<t_{1}^{gt}<t_{2}^{gt}<\dots$. These two sequences form a bipartite graph $\mathcal{G}(S_{p},S_{gt})$. The prediction is connected to the ground truth event if their types are equal. If $M(S_{p},S_{gt})$ is a set of all possible matches in the graph $\mathcal{G}(S_{p},S_{gt})$, OTD is computed as follows:
@@
where $U_{p}(m)$ is the number of unmatched predictions in matching $m$, $U_{gt}(m)$ is the number of unmatched ground truth events, $C_{ins}$ is an insertion cost, and $C_{del}$ is a deletion cost. It is common to take $C_{ins}=C_{del}$ [mei2019imputing]. It has been proven that OTD is a valid metric, as it is symmetric, equals zero for identical sequences, and satisfies the triangle inequality. However, OTD has practical limitations, which are described in the following section.
3 Limitations of the Next-Event and OTD Metrics
Figure 2: Comparison of OTD and T-mAP metrics. In example 1.a, prefix evaluation of three events using OTD results in an incorrect false positive. T-mAP addresses this issue by comparing all events within the time horizon, as shown in example 1.b. Additionally, OTD evaluates only the label with the maximum probability. From the OTD perspective, cases 2.a and 2.b have the same quality. In contrast, T-mAP evaluates the entire distribution, yielding a low score of 0.33 in case 2.a and a high score of 1 in case 2.b.
A predominant evaluation metric in the MTPP domain is the quality of next-event prediction. This metric typically comprises two distinct components: the accuracy (or error rate) of the next label prediction and either Mean Absolute Error (MAE) or Root Mean Square Error (RMSE) for time step prediction. However, these metrics do not assess the model’s ability to predict multiple future events. For instance, autoregression uses its own outputs as inputs for the next step, which can potentially lead to cumulative prediction errors. Consequently, long-horizon evaluation metrics are necessary.
The OTD metric extends evaluation to sequences by comparing fixed-size prefixes, but this approach limits its flexibility in assessing models with imprecise time step predictions, as illustrated in Fig. 2.1. If the model predicts events too frequently, the first $K$ predicted events will correspond to the early part of the ground truth sequence, resulting in false negatives. This outcome is inaccurate because allowing the model to predict more events would alter the evaluation. Conversely, if the time step is too large, the first $K$ predicted events will extend far beyond the horizon of the first $K$ target events, leading to a significant number of false positives. This is undesirable as it prevents the inclusion of additional ground truth events in the evaluation. Therefore, evaluating a dynamic number of events is essential to properly align with the time horizon in the ground truth.
Another limitation of the OTD metric is its inflexibility in evaluating predicted distributions of labels, as shown in Fig. 2.2. OTD evaluates only the labels with the highest probability, disregarding the full distribution. As a result, it relies on the calibration of model outputs and cannot assess the performance beyond a limited set of popular labels. However, models typically predict probabilities for all classes, allowing for an independent assessment of prediction quality for each label. Therefore, we aim to develop an evaluation metric that accurately reflects the performance across common and rare classes.
@@ 4 Temporal mAP: Prediction as Detection
In this section, we introduce a novel metric, Temporal mAP (T-mAP), inspired by object detection methods from computer vision [everingham2010pascalcvdetection], as illustrated in Fig. 1. The goal of object detection is to localize objects within an image and identify their types. In event sequences, we tackle a similar problem, but we consider the time dimension instead of horizontal and vertical axes. Unlike object detection, where objects have spatial size, each event in an MTTP is a point without a duration. Therefore, we replace the intersection-over-union (IoU) similarity between bounding boxes with the absolute time difference.
@@ 4.1 Definition
A prediction $i$ can be matched with ground truth $j$ iff $|t^{p}_{i}-t^{gt}_{j}|<\delta$. T-mAP identifies the matching that maximizes precision and recall, matching with maximum cover $c$. The precision of this matching is calculated as $c/|S^{l}_{>}(h)|$ and the recall as $c/|S^{l}_{gt}|$. For any threshold $h$, we can count true positives (TP), false positives (FP), and false negatives (FN) across all predicted and ground truth sequences. Note that there are no true negatives, as the model cannot explicitly predict the absence of an event. Similarly to binary classification, we can define a precision-recall curve by varying the threshold $h$ and estimating Average Precision (AP), the area under the precision-recall curve. T-mAP is the mean AP across all possible event types $l$.
@@ 4.2 Computation
Finding the optimal matching independently for each threshold value $h$ is impractical; thus, we need a more efficient method to evaluate T-mAP. In this section, we show how to optimize T-mAP computation using an assignment problem solver, like the Jonker-Volgenant algorithm [jonker1988linsum]. The resulting complexity of T-mAP computation is $\mathcal{O}(BN^{3})$, where $B$ is the number of evaluated sequences and $N=\max(n_{p},n_{gt})$ is the number of events within the horizon.
@@
For each pair of sequences $S^{l}_{p}$ and $S^{l}_{gt}$ we define a weighted bipartite graph $\mathcal{G}(S^{l}_{p},S^{l}_{gt})$ with $|S^{l}_{p}|$ vertices in the first and $|S^{l}_{gt}|$ vertices in the second part. For each pair of prediction $i$ and ground truth event $j$ with $|t^{p}_{i}-t^{gt}_{j}|<\delta$ we add an edge with weight $-s^{p}_{i}$, as shown in Fig. 3.a. Jonker-Volgenant algorithm finds the matching with the maximum number of edges in the graph, such that the resulting matching minimizes the total cost of selected edges, as shown in Fig. 3.b. We call this matching optimal matching and denote a set of all optimal matching possibilities as $M(\mathcal{G})$. For any threshold $h$, there is a subgraph $\mathcal{G}_{h}(S^{l}_{>}(h),S^{l}_{gt})$ with events whose scores are greater than $h$. The following theorem holds:
@@
According to this theorem, we can compute the matching for the prediction $S^{l}_{gt}$ and subsequently reuse it for all thresholds $h$ and subsequences $S^{l}_{>}(h)$ to construct a precision-recall curve for the entire dataset, as shown in Fig. 3.c. The proof of the theorem and the complete algorithm for T-mAP evaluation are provided in Appendix B.

### v2

@@ Abstract
Accurately forecasting multiple future events within a given time horizon is crucial for finance, retail, social networks, and healthcare applications. Event timing and labels are typically modeled using Marked Temporal Point Processes (MTPP), with evaluations often focused on next-event prediction quality. While some studies have extended evaluations to a fixed number of future events, we demonstrate that this approach leads to inaccuracies in handling false positives and false negatives. To address these issues, we propose a novel evaluation method inspired by object detection techniques from computer vision. Specifically, we introduce Temporal mean Average Precision (T-mAP), a temporal variant of mAP, which overcomes the limitations of existing long-horizon evaluation metrics. Our extensive experiments demonstrate that models with strong next-event prediction accuracy can yield poor long-horizon forecasts and vice versa, indicating that specialized methods are needed for each task. To support further research, we release HoTPP [^1] https://github.com/ivan-chai/hotpp-benchmark, the first benchmark designed explicitly for evaluating long-horizon MTPP predictions. HoTPP includes large-scale datasets with up to 43 million events and provides optimized procedures for both autoregressive and parallel inference, paving the way for future advancements in the field.
@@ 1 Introduction
The world is full of events. Internet activity, e-commerce transactions, retail operations, clinical visits, and numerous other aspects of our lives generate vast amounts of data in the form of timestamps and related information. In the era of AI, it is crucial to develop methods capable of handling these complex data streams. We refer to this type of data as Event Sequences (ESs). Event sequences differ fundamentally from other data types. Unlike tabular data (Wang & Sun, 2022), ESs include timestamps and possess an inherent order. In contrast to time series data (Lim & Zohren, 2021), ESs are characterized by irregular time intervals and structured descriptions of each event. These differences necessitate the development of specialized models and evaluation practices.
Sequence modeling is the primary task in Event Sequences (ESs) domain. In its simplest form, each event is defined by its type and occurrence time, a framework commonly known as Marked Temporal Point Processes (MTPP) (Rizoiu et al., 2017). Some studies extend MTPP to incorporate additional data fields and model complex dependencies between them (McDermott et al., 2024). However, the majority of MTPP approaches, as well as their evaluation pipelines, primarily focus on predicting the next event based on historical data.
In practice, a common question arises: what events will occur, and when, within a specific time horizon? Forecasting multiple future events presents unique challenges that differ from traditional next-event prediction tasks. For instance, autoregressive event sequence prediction involves applying the model to its own potentially erroneous predictions. However, the challenges of autoregressive prediction in the context of MTPP have not been thoroughly explored. Another difficulty lies in evaluation. Methods like Dynamic Time Warping (DTW) are typically unsuitable for event sequences due to strict ordering constraints (Su & Hua, 2017). Some works have applied Optimal Transport Distance (OTD), a variant of the Wasserstein distance, to compare sequences of predefined lengths. Still, the limitations of this metric have not been previously considered.
In this work, we provide the first in-depth analysis of models and metrics for long-horizon event forecasting, establishing a rigorous evaluation framework and a baseline for MTPP studies. Our key contributions are as follows:
@@
    We demonstrate that widely used evaluation methods for MTPPs often overlook critical aspects of model performance. We demonstrate that simple rule-based baselines can sometimes outperform popular deep learning methods when evaluated using OTD.
@@
    We introduce Temporal mean Average Precision (T-mAP), a novel evaluation metric inspired by best practices in computer vision. T-mAP evaluates variable length sequences within a specified time horizon, as illustrated in Figure 1. Unlike previous approaches, T-mAP accurately accounts for false positives and false negatives while being invariant to linear calibration. Additionally, we address a theoretical gap in computer vision by proving the correctness of the T-mAP computation algorithm.
@@
    Using our established methodology, we demonstrate that high next-event prediction accuracy does not necessarily translate into high-quality long-horizon forecasts; in many cases, our experiments show the opposite. This highlights the necessity of developing specialized models for the long-horizon prediction task.
@@
    We release HoTPP, a new open-source benchmark designed to facilitate long-horizon event sequence prediction research. HoTPP brings together datasets and methods from various domains, including financial transactions, social networks, healthcare, and recommender systems, greatly expanding the diversity and scale of data compared to prior benchmarks. Additionally, we offer an efficient inference algorithm necessary for large-scale evaluations.
@@
Figure 1: Temporal mAP (T-mAP) evaluation pipeline. Unlike previous methods, T-mAP evaluates sequences of variable lengths within the prediction horizon. It further improves performance assessment by analyzing label distributions rather than relying on fixed predictions.
2 Related Work
@@
Marked Temporal Point Processes. A Marked Temporal Point Process (MTPP) is a stochastic process consisting of a sequence of pairs $(t_{1},l_{1}),(t_{2},l_{2}),\dots$, where $t_{1}<t_{2}<\dots$ represent event times and $l_{i}\in\{1,\dots,L\}$ denote event type labels (Rizoiu et al., 2017). Common approaches to MTPP modeling focus on predicting the next event. A basic solution involves independently predicting the event time and type. A more advanced approach splits the original sequence into subsequences, one for each event type, and independently models each subsequence’s timing. Depending on the time-step distribution, the process is called Poisson or Hawkes.
Over the last decade, the focus has shifted toward increasing model flexibility by applying neural architectures. Several works have employed classical RNNs (Du et al., 2016; Xiao et al., 2017; Omi et al., 2019) and transformers (Zuo et al., 2020; Zhang et al., 2020), while others have proposed architectures with continuous time (Mei & Eisner, 2017; Rubanova et al., 2019; Kuleshov et al., 2024). In this work, we evaluate MTPP models with both discrete and continuous time architectures. Unlike previous benchmarks, we also assess simple rule-based predictors and popular methods from related fields, including GPT-like prediction models for event sequences (McDermott et al., 2024; Padhi et al., 2021) and Next-K models from time series analysis (Lim & Zohren, 2021). For more details on MTPP modeling, refer to Appendix A.
MTPP Evaluation. Previous benchmarks have primarily focused on medical data or traditional MTPP datasets. EventStream-GPT (McDermott et al., 2024) and TemporAI (Saveliev & van der Schaar, 2023) consider only medical data and do not implement methods from the MTPP field, despite their applicability. Early MTPP benchmarks, such as Tick (Bacry et al., 2017) and PyHawkes [^2] https://github.com/slinderman/pyhawkes, implement classical machine learning approaches but exclude modern neural networks. While PoPPy (Xu, 2018) and EasyTPP (Xue et al., 2023) include neural methods, they do not consider rule-based and Next-K approaches. Furthermore, PoPPy does not evaluate long-horizon predictions at all. EasyTPP addresses this limitation to some extent by reporting the OTD metric, though it does not provide the corresponding evaluation code.
Previous work also highlighted the limitations of Dynamic Time Warping (DTW) for event sequence evaluation (Su & Hua, 2017), showing that DTW’s strict ordering constraints are impractical and should be avoided. This work proposed an alternative, the Order-preserving Wasserstein Distance (OPW). In our study, we address the limitations of OTD, a variant of the Wasserstein Distance, and propose a new metric called T-mAP. Unlike OPW, T-mAP operates on timestamps rather than event indices and involves two hyperparameters, compared to three in OPW.
3 Limitations of the Next-Event and OTD Metrics
Figure 2: Comparison of OTD and T-mAP metrics. In example 1.a prefix evaluation of three events using OTD results in an incorrect, false negative. T-mAP addresses this issue by comparing all events within the time horizon, as shown in example 1.b. Additionally, OTD evaluates only the label with the maximum probability. From the OTD perspective, cases 2.a and 2.b have the same quality. In contrast, T-mAP evaluates the entire distribution, yielding a low score of 0.33 in case 2.a and a high score of 1 in case 2.b.
MTPPs are typically evaluated based on the accuracy of next-event predictions, with time and type predictions assessed independently. The quality of type predictions is measured by the error rate, while time prediction error is evaluated using either Mean Absolute Error (MAE) or Root Mean Squared Error (RMSE). However, these metrics do not account for the model’s ability to predict multiple future events. For example, in autoregressive models, the outputs are fed back as inputs for subsequent steps, which can lead to cumulative prediction errors. As a result, long-horizon evaluation metrics are necessary.
@@ 3 Limitations of the Next-Event and OTD Metrics
Recent studies have advanced the measurement of horizon prediction quality by employing Optimal Transport Distance (OTD) (Mei et al., 2019). This metric is analogous to the edit distance between the predicted event sequence and the ground truth. Suppose there is a predicted sequence $S_{p}=\{(t^{p}_{i},l^{p}_{i})\}_{i=1}^{n_{p}},0<t_{1}^{p}<t_{2}^{p}<\dots$, and a ground truth sequence $S_{gt}=\{(t^{gt}_{j},l^{gt}_{j})\}_{j=1}^{n_{gt}},0<t_{1}^{gt}<t_{2}^{gt}<\dots$. These two sequences form a bipartite graph $\mathcal{G}(S_{p},S_{gt})$. The prediction is connected to the ground truth event if their types are equal. Denote $M(S_{p},S_{gt})$ a set of all possible matches in the graph $\mathcal{G}(S_{p},S_{gt})$. OTD finds the minimum cost among all possible matchings:
@@
where $U_{p}(m)$ is the number of unmatched predictions in matching $m$, $U_{gt}(m)$ is the number of unmatched ground truth events, $C_{ins}$ is an insertion cost, and $C_{del}$ is a deletion cost. It is common to take $C_{ins}=C_{del}$ (Mei et al., 2019). It has been proven that OTD is a valid metric, as it is symmetric, equals zero for identical sequences, and satisfies the triangle inequality.
OTD is computed between fixed-size prefixes, but this approach limits metric flexibility in assessing models with imprecise time step predictions, as illustrated in Fig. 2.1. If the model predicts events too frequently, the first $K$ predicted events will correspond to the early part of the ground truth sequence, resulting in false negatives. This outcome is inaccurate because allowing the model to predict more events would alter the evaluation. Conversely, if the time step is too large, the first $K$ predicted events will extend far beyond the horizon of the first $K$ target events, leading to a significant number of false positives. This is undesirable as it prevents the inclusion of additional ground truth events in the evaluation. Therefore, evaluating a dynamic number of events is essential to properly align with the ground truth’s time horizon.
MTPP evaluation metrics can be categorized into two groups: those considering event indices and those assessing time and label prediction quality independently of actual positions. Next-event metrics rely on ordering, even when events share identical timestamps, making the order meaningless. While Optimal Transport Distance (OTD) itself is invariant to order, the extraction of sequences for comparison is influenced by event indices. For example, when OTD compares prefixes of length $K$, the algorithm must place $K$ correctly predicted events at the beginning of a sequence to minimize the OTD score. As a result, OTD, like next-event metrics, can depend on event ordering even when the order cannot be uniquely determined. This dependency is non-trivial and difficult to measure or control accurately.
Another limitation of the OTD metric is its inability to evaluate the full predicted distribution of labels, as shown in Figure 2.2. OTD considers only the labels with the highest probability, ignoring the complete distribution. This makes it dependent on model calibration and limits the ability to assess performance across a broader range of event types, such as long-tail predictions. However, models typically predict probabilities for all classes, allowing for a more comprehensive assessment. Therefore, we aim to develop an evaluation metric that accurately captures performance across common and rare classes.
@@ 4 Temporal mAP: Prediction as Detection
In this section, we introduce a novel metric, Temporal mAP (T-mAP), which analyses all errors within a predefined time horizon, explicitly controls ordering, and is invariant to linear calibration. T-mAP is inspired by object detection methods from computer vision (Everingham et al., 2010), as illustrated in Fig. 1. The goal of object detection is to localize objects within an image and identify their types. In event sequences, we tackle a similar problem but consider the time dimension instead of horizontal and vertical axes. Unlike object detection, where objects have spatial size, each event in an MTTP is a point without a duration. Therefore, we replace the intersection-over-union (IoU) similarity between bounding boxes with the absolute time difference.
@@ 4.1 Definition
By definition, a predicted event $i$ can be matched with a ground truth event $j$ iff $|t^{p}_{i}-t^{gt}_{j}|\leq\delta$, meaning the time difference between the predicted and ground truth events is less than or equal to $\delta$. T-mAP identifies the matching that maximizes precision and recall, i.e., matching with maximum cover $c$. The precision of this matching is calculated as $c/|S^{l}_{>}(h)|$ and the recall as $c/|S^{l}_{gt}|$. For any threshold $h$, we can count true positives (TP), false positives (FP), and false negatives (FN) across all predicted and ground truth sequences. Note that there are no true negatives, as the model cannot explicitly predict the absence of an event. Similar to binary classification, we can define a precision-recall curve by varying the threshold $h$ and then estimate the Average Precision (AP), the area under the precision-recall curve. Finally, T-mAP is defined as the average AP across all classes.
@@ 4.2 Computation
Finding the optimal matching independently for each threshold value $h$ is impractical; thus, we need a more efficient method to evaluate T-mAP. This section shows how to optimize T-mAP computation using an assignment problem solver, like the Jonker-Volgenant algorithm (Jonker & Volgenant, 1988). The resulting complexity of T-mAP computation is $\mathcal{O}(BN^{3})$, where $B$ is the number of evaluated sequences and $N=\max(n_{p},n_{gt})$ is the number of events within the horizon.
@@
For each pair of sequences $S^{l}_{p}$ and $S^{l}_{gt}$ we define a weighted bipartite graph $\mathcal{G}(S^{l}_{p},S^{l}_{gt})$ with $|S^{l}_{p}|$ vertices in the first and $|S^{l}_{gt}|$ vertices in the second part. For each pair of prediction $i$ and ground truth event $j$ with $|t^{p}_{i}-t^{gt}_{j}|\leq\delta$ we add an edge with weight $-s^{p}_{i}$, equal to negative logit of the target class, as shown in Fig. 3.a. Jonker-Volgenant algorithm finds the matching with the maximum number of edges in the graph, such that the resulting matching minimizes the total cost of selected edges, as shown in Fig. 3.b. We call this matching optimal matching and denote a set of all optimal matching possibilities as $M(\mathcal{G})$. For any threshold $h$, there is a subgraph $\mathcal{G}_{h}(S^{l}_{>}(h),S^{l}_{gt})$ with events whose scores are greater than $h$. The following theorem holds:
@@
According to this theorem, we can compute the matching for the prediction $S^{l}_{gt}$ and subsequently reuse it for all thresholds $h$ and subsequences $S^{l}_{>}(h)$ to construct a precision-recall curve for the entire dataset, as shown in Fig. 3.c. The proof of the theorem, the study of calibration dependency, and the complete algorithm for T-mAP evaluation are provided in Appendix B.
4.3 T-mAP Hyperparameters
T-mAP has two hyperparameters: the maximum allowed time delta $\delta$ and evaluation horizon $H$. We set $\delta$ twice the cost of the OTD because when the model predicts an incorrect label, OTD removes the prediction and adds the ground truth event with the total cost equal to $2C$. The horizon $H$ must be larger than $\delta$ to evaluate timestamp quality adequately. Therefore, depending on the dataset, we select $H$ to be approximately 3-7 times larger than $\delta$, ensuring the horizon captures an average of 6-15 events. The empirical study of T-mAP hyperparameters is presented in Figure 4. As shown, the chosen $\delta$ parameter is positioned to the right of the initial slope of the parameter-quality curve, ensuring an optimal balance between time granularity and task difficulty.
Figure 4: T-mAP dependency on the $\delta$ parameter. The dashed line indicates the selected value. Results for the StackOverflow and Amazon datasets are provided in the Appendix E.